OpenStudy (anonymous):

what are the coordinates of the points 3 units from the y-axis and at distance squareroot of 5 from (5,3)?

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

So this time, with the distance formula \[d =\sqrt{ (x_2-x_1)^2+(y_2-y_1)^2} \] you have one set of ordered pairs \[ (5,3)=(x_1,y_1)\] and a value for d \[d=\sqrt {5}\] What do you know about the other points?

OpenStudy (lucaz):

3 units from y-axis = (3,y) ?

OpenStudy (anonymous):

almost - (±3,y)

OpenStudy (anonymous):

i dont knw that sample part of my reviewer

OpenStudy (anonymous):

it doesn't say which direction the units are (that's the tricky part)

OpenStudy (lucaz):

right

OpenStudy (anonymous):

huh can you help me

OpenStudy (anonymous):

So, @jacalneaila , It's like the other problem, except you're solving for y given a distance \[\sqrt{5} =\sqrt{ (3-5)^2+(y-3)^2} \]

OpenStudy (anonymous):

You know that the first part of the ordered pair (the x component) is either (+3,y) or (-3,y), so you solve for 2 solutions of y

OpenStudy (anonymous):

\[\sqrt{5} =\sqrt{ (3-5)^2+(y-3)^2} \] \[5 =(3-5)^2+(y-3)^2 \] \[5 =4+y^2-6y+9 \] \[y_2-6y+8=0\] can you factor that?

OpenStudy (anonymous):

factor?

OpenStudy (anonymous):

what next

OpenStudy (anonymous):

\[y^2-6y+8=0\] \[(y-4)(y-2)=0\] so at (+3,y) there are two values of y that are a distance of sqrt 5 from point (5,3) y=4 y=2 So the two points would be (3,4) and (3,2) Now we have to do the same thing for (-3,y) \[\sqrt 5 =\sqrt{ (-3-5)^2+(y-3)^2}\] \[5=(8)^2+(y^2-6y+9)\] \[5=64+y^2-6y+9\] \[y^2-6y-68=0\] then the roots are \[y=\frac{6±\sqrt{36+272}}2\] \[y=3±\sqrt {77}\] although I dunno if that's right

OpenStudy (anonymous):

huh.. i will review that

OpenStudy (anonymous):

next example

OpenStudy (anonymous):

How many examples do you have?

OpenStudy (anonymous):

in my reivewer its 50 examples .. i have exam tommorow its 200 question

OpenStudy (anonymous):

I can help you with a few more I think ^^

OpenStudy (anonymous):

I don't think I'll be able to last through all 50 though!

OpenStudy (anonymous):

Just pick the one or two that you're having the most trouble with, unless you think that they're really short, then maybe a few more than 2.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

can close this? then next sample

OpenStudy (anonymous):

sure